OpenStudy (anonymous):

can someone please help me!! y varies inversely with x: if x=2 and y=25 what is x when y is 5?

OpenStudy (anonymous):

y varies inversely with let y=k/x put x=2 and y=25 25=k/2 k=50 therefore, y=50/x put y=5 5=50/x x=10

OpenStudy (anonymous):

thankyou!:)